# PINN-MEP: Continuous Neural Representations for Minimum-Energy Path Discovery in Molecular Systems

### v5

@@ 3 Experiments
We evaluate our method on two molecular systems: the AIB9 peptide and Bovine Pancreatic Trypsin Inhibitor (BPTI). These systems represent different scales of complexity and serve to validate our method’s ability to discover physically meaningful transition paths. In Appendix A.2, we compare and benchmark our method with the related approaches outlined in (Du et al., 2024; Lee et al., 2025). In Appendix B.1, we present ablation studies analyzing the impact of different embedding strategies, interpolation methods, neural network architectures, and regularization approaches.
@@ 3.1 AIB9 Transition Path Discovery
We first examine the AIB9 system, a 9-residue artificial protein with 129 atoms that exhibits two well-defined metastable states, making it an ideal test case for transition path methods. This system effectively bridges abstract models and complex proteins: it is small enough for extensive reference simulations yet complex enough to exhibit realistic conformational changes; its energy landscape is well-characterized; and its transitions are easily visualized through dihedral angle projections. Unlike artificial toy problems, AIB9 allows direct validation with actual biomolecular force fields. We use the AMBER ff15ipq-m force field for protein mimetics (Bogetti et al., 2020) implemented in DMFF (Wang et al., 2023), and we generate multiple MEPs between the system’s two metastable states.
Figure 2 shows the discovered transition paths projected onto the central residue’s - dihedral angle space. The method identifies multiple physically plausible pathways between the states, which are also observed with long MD simulations. For each path, different conformations of the states were picked as start and end points; this, along with the neural network’s random initialization, leads to the discovery of multiple different transitions. Notably, the discovered MEPs align well with the free energy landscape despite our choice of action functional without the trajectory entropy term and the assumption of overdamped dynamics, which may not be ideal for this vacuum system, but is less problematic for hydrated systems.
@@ 4 BPTI Conformational Change Pathway
Bovine Pancreatic Trypsin Inhibitor (BPTI) is a clinically significant protein used as an anticoagulant in medical procedures. BPTI is 58 residues long (892 atoms) and has complex conformational changes involving disulfide bond rearrangements paired with backbone changes.
To investigate BPTI’s conformational transitions, we use as reference structures five key states from a landmark molecular dynamics (MD) simulation performed by D.E. Shaw Research on their Anton supercomputer (Shaw et al., 2010). These publicly available snapshots provide rare benchmark structures for validating our approach to BPTI’s transition mechanisms, making it an excellent larger system to test the method. Full details of our BPTI system preparation, explicit solvent shell construction, and hybrid neural network architecture for handling water molecules are provided in Appendix B.2.
@@
We matched their simulation conditions by incorporating explicit water molecules in our model to ensure a meaningful comparison with the Anton simulation. We generated water shells extending 14Å from any protein atom for both the initial and final BPTI conformations. This increases the number of atoms from 892 to 8,300. To handle water molecules during transitions, we pair proximal water molecules between start and end states using the Hungarian algorithm (Kuhn, 1955). We then model their movement using simple parameterized splines, rather than directly incorporating all water coordinates into the neural network. This hybrid neural-spline approach allows us to include explicit solvent effects while keeping computational costs low. The specific details of our water shell construction and matching algorithm are described in Appendix B.2.
@@
While our approach does not provide the same statistical ensemble information as equilibrium MD, it efficiently captures the key transition pathway, which is often the primary mechanistic insight sought by MD practitioners. For most scientifically interesting systems—typically large biomolecular complexes—equilibrium sampling of the complete conformational space or transition path ensemble remains computationally prohibitive, making identification of dominant pathways particularly valuable. It’s worth noting that even the reference millisecond-scale simulation may not have fully achieved equilibrium sampling of the conformational space. Appendix B.2.7 provides a detailed computational performance comparison.
@@ 5 Limitations
While our feedforward neural architecture can model conformational transitions, it could benefit from incorporating domain-specific structural priors to improve training efficiency and convergence speed. Furthermore, physical force fields introduce challenges through high-energy spikes that can cause training instability and divergence in some runs. Our gradient clipping approach, outlined in B.2.5, largely mitigates this issue but doesn’t universally prevent training failures.
The method has been validated primarily on small to medium-sized protein systems, and scaling to larger multi-domain proteins could require additional optimization strategies. Our approach is also constrained to systems with suitable differentiable force fields, but this limitation diminishes as more implementations become available.
Finally, while computationally efficient, our method focuses on identifying MEPs rather than sampling the full statistical ensemble of transition pathways, representing a trade-off between efficiency and exhaustively sampling a system. However, as noted in the BPTI section 4, this mechanistic insight is often a key objective for MD practitioners studying complex biomolecular transitions.
@@ 6 Conclusion
We demonstrated that reformulating MEP discovery as a continuous neural optimization problem enables efficient discovery of molecular transition mechanisms with unprecedented computational efficiency. Our approach transforms the complex Onsager-Machlup action functional into a straightforward energy-based loss function that eliminates the need for force and path gradient computations required by related methods, which are known to induce optimization challenges in PINNs. This work provides a practical tool for molecular dynamics practitioners that captures essential mechanistic insights without requiring specialized hardware or extensive simulation time.
While our current implementation shows promising results, several concrete directions for improvement remain. Future work should explore neural architectures that incorporate domain-specific structural priors beyond simple MLPs with sinusoidal/Fourier embeddings. Related work by Ramakrishnan et al. (2025) investigates non-uniform sampling schemes along the path that might further aid convergence for paths that constitute dramatic rearrangements. Integrating pretrained generative models suggested by Raja et al. (2025) could provide better initial transition guesses, potentially accelerating convergence by starting optimization from conformations that already respect molecular geometry. Extending the method to handle larger multi-domain proteins is also a frontier of great scientific interest, which might bring new challenges. Future work could also explore ways to estimate free energies and reaction rates from our discovered paths.
7 Broader Impacts
Our research democratizes access to the study of larger molecular systems by enabling transition path studies on standard hardware. This may accelerate drug discovery and protein engineering while contributing to sustainable computing through drastically reduced energy consumption. We see no direct negative societal impacts, though we acknowledge that accelerated protein engineering capabilities could be misused in the future, and we encourage responsible use with appropriate experimental validation.
Appendix A Technical Appendices and Supplementary Material
@@ A.1 Extended Derivation of the energy-based Loss Function
We use a first-order Taylor expansion to relate energy differences to gradient magnitudes
@@
For a MEP, the displacement tends to align either parallel or anti-parallel with the gradient direction (i.e., for path segments)
@@
While the last two approximations do not change the location of the global minimum, they do affect the optimization landscape before that minimum is reached.
@@ A.1.1 From the Onsager-Machlup Action to a geometric Line Integral
with equality when is parallel or anti-parallel to , with the direction determined by whether the path is ascending or descending the energy landscape. This is precisely the defining characteristic of the MEP - a path that follows the potential energy landscape’s gradient while avoiding high-energy regions.
@@ A.1.2 Discretization and Connection to Energy Differences
where is the remainder of the order and is the angle between the gradient and displacement vectors.
@@
This forms a surrogate objective with the same minimum as our original action.
@@ A.1.4 Neural Network Implementation
By minimizing through gradient-based optimization, the neural network learns to represent low-energy pathways connecting the stable states, effectively discovering the MEP predicted by the Onsager-Machlup framework. The resulting path follows the potential energy landscape’s gradient while avoiding high-energy regions, representing the most probable transition mechanism in the overdamped limit.
A.2 Connection between Doob’s Lagrangian and MEP Optimization

### v6

@@ 3 Experiments
We evaluate our method on two molecular systems: the AIB9 peptide and the Bovine Pancreatic Trypsin Inhibitor (BPTI). These systems represent different scales of complexity and serve to validate our method’s ability to discover physically meaningful transition paths. In Appendix B, C and D, we show extensive benchmarks of our method with the related approaches outlined in (Du et al., 2024; Lee et al., 2025) and classical MEP approaches, namely chain-of-states, NEB, and the string method, and demonstrate that we outperform all other methods in terms of both mean and maximum path energy. In Appendix F, we present ablation studies analyzing the impact of our architectural choices on MEP quality.
@@ 3.1 AIB9 Transition Path Discovery
We first examine the AIB9 system, a 9-residue artificial protein with 129 atoms that exhibits two well-defined metastable states, making it an ideal test case for transition path methods. It is small enough for extensive reference simulations yet complex enough to exhibit realistic conformational changes, and its transitions are easily visualized through dihedral angle projections. We therefore also use it for the majority of benchmarking and ablation experiments. Unlike artificial toy problems, AIB9 allows direct validation with actual biomolecular force fields. We use the AMBER ff15ipq-m force field for protein-mimetics (Bogetti et al., 2020) implemented in DMFF (Wang et al., 2023), and we generate multiple MEPs between the system’s two metastable states.
Figure 2 shows the discovered transition paths projected onto the central residue’s - dihedral angle space. The method identifies multiple physically plausible pathways between the states, which are also observed with long MD simulations. For each path, different conformations in the states were picked as start and end points; this, along with the neural network’s random initialization, leads to the discovery of multiple different transitions. Notably, the discovered MEPs align well with the free energy landscape despite our choice of an action functional without the trajectory entropy term and the assumption of overdamped dynamics, which may not be ideal for this vacuum system, but is less problematic for hydrated systems.
@@ 4 BPTI Conformational Change Pathway
Bovine Pancreatic Trypsin Inhibitor (BPTI) is a clinically significant protein used as an anticoagulant in medical procedures. BPTI is 58 residues long (892 atoms) and exhibits complex conformational changes involving disulfide bond rearrangements paired with backbone changes.
To investigate BPTI’s conformational transitions, we use as reference structures five states from a landmark molecular dynamics (MD) simulation performed by D.E. Shaw Research on their Anton supercomputer (Shaw et al., 2010). These publicly available snapshots provide rare benchmark structures for validating our approach to BPTI’s transition mechanisms, making it an excellent system to test the method on a larger scale. Full details of our BPTI system preparation and hydration are provided in Appendix G.
@@
We matched their simulation conditions by incorporating explicit water molecules to ensure a meaningful comparison with the Anton simulation. We hydrate both the initial and final BPTI conformations, as done in a conventional MD system setup. This increases the number of atoms from 892 to 3,500. To handle water molecules during transitions, we pair proximal water molecules between start and end states using the Hungarian algorithm (Kuhn, 1955).
@@
While our approach does not provide the same statistical ensemble information as the reference MD, it efficiently captures the key transition pathway. It is worth noting that even the reference millisecond-scale simulation has not achieved equilibrium sampling, as it is comprised of a single transition. Appendix I provides a detailed computational performance comparison.
5 Comparative Evaluation and Ablation Studies
To systematically validate our approach, we conducted extensive benchmarks against established methods and ablation studies examining our architectural choices across three molecular systems of increasing complexity. We evaluate path quality using two key metrics: the maximum energy along the path (indicating transition state barriers), the arc-length adjusted mean energy (accounting for how quickly paths traverse high-energy regions), and the standard deviation of this quantity for non-deterministic methods. The full results for the benchmarks are presented in sections B, C, and D, and the ablations are discussed in section F of the appendix, respectively.
Benchmark Results: AdaPath consistently outperforms classical approaches (Chain-of-States, String Method, NEB) and deep learning approaches like Doob’s Lagrangian in both computational efficiency and path quality. For the alanine dipeptide, we achieve superior energy barriers (-86.6 vs. -71.1 kJ/mol for Chain-of-States) while requiring fewer force evaluations (6.4e4 vs. 2.4e5). Here and on other systems, Chain-of-States emerged as the best-performing classical method. We note that Doob’s Lagrangian targets the more ambitious objective of modeling the entire transition path ensemble rather than discovering a single MEP, making direct performance comparisons inherently limited.
This performance advantage scales to larger systems: on AIB9 (129 atoms), AdaPath reaches -644 kJ/mol maximum energy compared to Chain-of-States’ -461 kJ/mol and substantially outperforms all other tested methods (Appendix C). On the BPTI system with explicit solvent (3,500 atoms), classical methods either fail to converge and produce erratic paths that bypass intermediate reference structures from the landmark D.E. Shaw simulation, as shown qualitatively in Appendix D.
Architectural Validation: Ablation studies confirm our core design choices (Appendix F). Parameter sharing proves essential for scalability; standard MLPs fail catastrophically on large systems, with even billion-parameter variants showing poor performance on BPTI. Our optimizer analysis reveals that Muon maintains stable path continuity, while Adam exhibits erratic ”teleportation” behavior that rapidly jumps between states without traversing transition regions in small systems. The logarithmic energy transformation proves critical for training stability, reducing energy variance by orders of magnitude.
@@ 6 Limitations
While our approach demonstrates strong performance on protein systems up to 3,500 atoms, several limitations should be noted. First, our method is constrained to systems with suitable differentiable force fields, though this limitation is diminishing as more implementations become available. Second, our method produces single representative MEPs rather than sampling the complete statistical ensemble of transition pathways. While this represents a trade-off between computational efficiency and exhaustive sampling, the mechanistic insights provided are often the primary objective for practitioners studying biomolecular transitions.
@@ 7 Conclusion
We demonstrated that reformulating MEP discovery as a continuous neural optimization problem, combined with a scalable neural architecture, enables efficient discovery of molecular transition mechanisms for large biomolecular systems in explicitly modeled solvent. Our approach makes two key contributions: deriving a simple energy-based loss function directly from the Onsager-Machlup action functional, and introducing AdaPath. This architecture scales favorably with system size via parameter sharing across the atom dimension. By encoding path continuity through architectural design rather than explicit loss terms, we sidestep known optimization challenges in PINNs and multi-loss training. Several promising directions for future development emerge from this work. The scaling properties of AdaPath suggest the method could be extended to even larger biomolecular systems, such as membrane proteins or multi-domain complexes that routinely contain tens of thousands of atoms in MD simulations, not counting water. Our discovered MEPs could serve as guidance paths for enhanced sampling methods, enabling free energy calculations through escorted Jarzynski-like estimators (Jarzynski, 1997; Vaikuntanathan & Jarzynski, 2011) by using the MEP as a basis for a non-equilibrium pulling protocol between stable states. The architecture and loss function may also prove valuable for other non-biological physical systems where rare events might be of interest.
Appendix A Appendix
@@ A.1 Extended Derivation of the energy-based Loss Function
We use a first-order Taylor expansion to relate energy differences to gradient magnitudes, which requires displacement vectors to be sufficiently small
@@
For a MEP, the displacement tends to align either parallel or antiparallel with the gradient direction (i.e., for path segments)
@@
While the last two approximations do not change the location of the global minimum, they do affect the optimization landscape before that minimum is reached. Critically, when the Taylor approximation breaks down due to large displacement vectors, the resulting energy-based loss can admit pathological solutions where paths remain in stable states and then rapidly ”teleport” to avoid high-energy transition regions.
@@ A.1.1 From the Onsager-Machlup Action to a geometric Line Integral
With equality when is parallel or antiparallel to , with the direction determined by whether the path is ascending or descending the energy landscape. This is precisely the defining characteristic of the MEP - a path that follows the potential energy landscape’s gradient while avoiding high-energy regions.
@@ A.1.2 Discretization and Connection to Energy Differences
where is the angle between the gradient and displacement vectors.
Critically, this Taylor expansion approximation is only valid when is sufficiently small for the higher-order term to be negligible. When displacement vectors become large, the approximation breaks down. It can lead to pathological solutions, where the optimization finds paths that avoid high-energy transition regions by rapidly ”teleporting” between stable states, rather than discovering smooth, minimum-energy routes.
@@
This forms a surrogate objective with the same minimum as our original action, provided the Taylor approximation remains valid throughout optimization.
@@ A.1.4 Neural Network Implementation
By minimizing through gradient-based optimization, the neural network learns to represent low-energy pathways connecting the stable states, effectively discovering the MEP predicted by the Onsager-Machlup framework. The resulting path follows the potential energy landscape’s gradient while avoiding high-energy regions, representing the most probable transition mechanism in the overdamped limit. However, as noted above, the validity of this approach depends critically on maintaining sufficiently small displacement vectors during optimization to ensure the underlying Taylor approximation remains valid.
Appendix B Benchmarks: Alanine Dipeptide
We provide a direct comparison between our AdaPath method, Du et al. (2024), Lee et al. (2025), as well as classical, well-established methods like the string method, chain-of-states, and nudged elastic band, in terms of computational efficiency, as evaluated by the number of force field/energy function calls, and the maximum energies observed along the paths, corresponding to the peak of the potential barrier, both as a mean across five seeds as well as the lowest value achieved across all runs. We further compute the arc-length-adjusted mean energy along the path. We chose this metric as naive averaging along a discretized path does not account for how fast a path is traversing a high-energy region; it could, therefore, shortcut a region of high energy but have a very low non-adjusted energy. To provide a rigorous comparison, we conducted benchmarks using experimental conditions identical to those reported by Du et al. for alanine dipeptide: AMBER14 force field, 300 K temperature, and also referred to their results when applicable, but adjusted the significant figure notation. Table 1 presents these results alongside those reported in Du et al. (2024) and their recent follow-up work Lee et al. (2025). The only exception is that we do not evaluate trajectory probability, as the assumption of optimization over a non-physical path length makes calculating this quantity nonsensical. Furthermore, the arc-length-adjusted energy quantities are not applicable to Doob’s method, as they do not model a single path but rather the entire transition path ensemble.
For the classical methods, we conducted extensive hyperparameter sweeps to ensure fair comparison and optimal performance. Chain-of-States was optimized across path discretizations (20-60 images), spring constants (1e-7 to 1e-1), and learning rates (5e-2 to 1e-1), with the best configuration using 40 images, a spring constant of 1e-6, and the Adam optimizer with a learning rate of 5e-2. The String Method was evaluated across image counts (20-60), timesteps (0.1-4000), and reparameterization frequencies (1-10 steps), achieving optimal performance with 40 images, timestep 3900, and reparameterization every three steps. NEB was tested across various path discretizations (20-60 images), spring constants (1e-7 to 1e-1), and learning rates (1e-6 to 1e-1), with the best results obtained using 60 images, a spring constant of 2e-5, and the Adam optimizer with a learning rate of 1e-5. These configurations represent the optimal settings identified through systematic exploration of the hyperparameter space for each method. All classical approaches we study are deterministic, and hence are only run once, unlike the stochastic approaches, which are run five times.
Since the field lacks standardized benchmarks, we believe that building on the work from (Du et al., 2024) could be a helpful step toward having a more standardized approach to methods aiming to solve for MEPs, the transition path ensemble, or other similar approaches. We want to reiterate that (Du et al., 2024) is more ambitious in its sampling objectives, and therefore, one-to-one comparisons are not entirely appropriate. However, the transition state energy, as represented by the Max Energy entry, should be consistent between MEP approaches and methods that aim to sample the entire transition ensemble.
| Method | # Evals | Max Energy | MinMax Energy | Mean Energy |
| MCMC variable length | 2.1e7 | N/A | ||
| MCMC* | 1.3e9 | N/A | ||
| MCMC variable length | 1.9e8 | N/A | ||
| MCMC | 1e10 | N/A | N/A | N/A |
| MCMC variable length | 1e10 | N/A | N/A | N/A |
| MCMC | 1e10 | N/A | N/A | N/A |
| Doob’s Cartesian | 3.8e7 | N/A | ||
| w/ 2 Mixtures | 5.1e7 | N/A | ||
| w/ 5 Mixtures | 5.1e7 | N/A | ||
| Doob’s Internal | 3.8e7 | N/A | ||
| w/ 2 Mixtures | 5.1e7 | N/A | ||
| w/ 5 Mixtures | 5.1e7 | N/A | ||
| Doob’s Seq2Seq | – | N/A | ||
| w/ Annealing | – | N/A | ||
| w/ Interpolation | – | N/A | ||
| MaxLL | – | N/A | ||
| w/ Annealing | – | N/A | ||
| w/ Interpolation | – | N/A | ||
| Chain-of-States | 2.4e5 | |||
| String Method | 3.3e6 | |||
| NEB | 7.7e6 | |||
| AdaPath | 6.4e4 |
AdaPath requires approximately fewer energy function evaluations (6.4e4 vs. 2.4e5) compared to the best classical method (Chain-of-States) while achieving superior energy barriers ( vs. kJ/mol for the minimum maximum energy). Compared to the Doob’s Lagrangian approach, AdaPath requires approximately fewer energy function evaluations (6.4e4 vs. 5.1e7) while achieving significantly lower maximum energy barriers. The classical methods demonstrate substantial variation in computational efficiency, with Chain-of-States providing the best computational performance among baseline approaches in terms of function evaluations. However, AdaPath still requires fewer evaluations while achieving better energy minimization. The lower number of required evaluations for AdaPath is likely due to the efficient neural representation compared to discrete path sampling approaches used by classical methods, and the large number of images that these methods need to function well.
Appendix C Benchmarks: AIB9
We conducted similar benchmarks on the AIB9 peptide system, a 129-atom artificial protein that exhibits well-defined metastable states and realistic conformational changes. This system serves as an intermediate complexity benchmark between simple dipeptides and larger protein systems. For the classical methods, we performed comprehensive hyperparameter optimization: Chain-of-States was evaluated across path discretizations (20 to 60 images), spring constants (1e-6 to 1e-1), and learning rates (1e-3 to 1e-1), achieving optimal performance with 40 images, spring constant 1e-5, and Adam optimizer with learning rate 1e-2. String Method was tested across image counts (20 to 60), timesteps (1 to 20), and reparameterization frequencies (1 to 15 steps), with best results using 60 images, timestep 10, and reparameterization every three steps. NEB was optimized across path discretizations (20 to 60 images), spring constants (1e-7 to 1e-1), and learning rates (1e-6 to 1e-1), yielding an optimal configuration with 60 images, spring constant 1e-3, and Adam optimizer with learning rate 1e-5.
For the Doob’s Lagrangian method, we adopted the settings reported for the Chignolin system in Du et al. (2024), as it represents a similar-sized protein system. Due to observed training instability, we added gradient clipping (1e-2) to stabilize optimization. We also found the original parameterization to be slightly under-parameterized for this system and increased the MLP width to 768 per layer across five layers. While the resulting energies appear high, they are consistent with the 3000 kJ/mol values reported for Chignolin in the original work, suggesting this energy scale is characteristic of the method’s behavior on protein systems of this size.
| Method | # Evals | Max Energy | MinMax Energy | Mean Energy |
| Chain-of-States | 2.9e6 | |||
| String Method | 3.1e6 | 1.3e4 | 1.3e4 | 3.2e3 |
| NEB | 4.1e6 | 2.7e3 | 2.7e3 | |
| Doob’s Lagrangian | 1.2e7 | 5.8e3 | N/A | |
| AdaPath | 4.7e6 |
The results demonstrate AdaPath’s superior performance compared to both classical methods and the Doob’s Lagrangian approach on this complex system. AdaPath achieved the best overall energy minimization, with a minimum maximum energy of kJ/mol, substantially outperforming Chain-of-States ( kJ/mol), String Method (1.3e4 kJ/mol), NEB (2.7e3 kJ/mol), and Doob’s Lagrangian (5.8e3 kJ/mol). The arc-length-adjusted mean energies further highlight AdaPath’s advantages, achieving kJ/mol. Notably, AdaPath demonstrated both superior energy minimization and significantly improved consistency across multiple seeds, with 3 out of 5 seeds achieving excellent results below kJ/mol, while Doob’s Lagrangian showed high variability with a standard deviation of 7.8e3 kJ/mol. The substantial energy differences between methods highlight the importance of the optimization objective and neural architecture design for effective transition path discovery.
Appendix D Benchmarks: BPTI
We conducted benchmarks on the BPTI (Bovine Pancreatic Trypsin Inhibitor) system to evaluate our AdaPath method against classical transition path discovery approaches on a significantly larger and more complex biomolecular system, BPTI. The system setup, including explicit solvation, ionic strength, and force field parameters, is detailed in Appendix G.
For the classical methods (Chain-of-States, String Method, and NEB), we employed the same hyperparameter configurations that proved optimal for the AIB9 system rather than conducting comprehensive sweeps, as the computational cost of extensive hyperparameter optimization on a system of this size would be prohibitive. Specifically, we used the following methods: Chain-of-States with 40 images, a spring constant of 1e-5, and the Adam optimizer with a learning rate of 1e-2; String Method with 60 images, a timestep of 10, and reparameterization every three steps; and NEB with 60 images, a spring constant of 1e-3, and the Adam optimizer with a learning rate of 1e-5.
For AdaPath, we used the training parameters specified in Appendix G, which is a more parameter-efficient variant of the model that performed best on the smaller AIB9 system.
| Method | # Evals | Max Energy | MinMax Energy | Mean Energy |
| Chain-of-States | 1.4e6 | |||
| String Method | 3.8e6 | 1.4e4 | 1.4e4 | 4.3e3 |
| NEB | 1.1e6 | 1.6e5 | 1.6e5 | 2.3e5 |
| AdaPath | 1.9e5 |
The results demonstrate AdaPath’s superior performance on this complex, explicitly solvated system. Figure 4 shows that among classical approaches, Chain-of-States produces the most reasonable pathway but fails to navigate through all intermediate reference structures from the D.E. Shaw simulation. The String Method fails to converge to a good solution, while NEB displays a jagged path, a known issue for this method in high-dimensional systems (Zhang et al., 2016). In contrast, AdaPath successfully discovers transition paths that pass through all intermediate reference structures, as demonstrated in Figure 3, while achieving significantly lower maximum energy barriers.
Appendix E Other related Methods
While our primary benchmark comparison focuses on the Doob’s Lagrangian method due to similarity in task and system setup, examining how our approach relates to other methods for transition path discovery/MEP generation is also informative. Here, we present results from two related approaches in the literature. However, we emphasize that direct numerical comparisons should be treated with caution due to differences in force fields, optimization objectives, and units used.
| Method | # Evaluations | Max Energy (eV) |
| NEB | 8.1e3 | |
| INR | 3.0e3 |
| Method | CVs | # Evaluations | Runtime / Path |
| MCMC (Two-Way Shooting) | No | 1e9 | 100 hours |
| Metadynamics | Yes | 1e6 | 10 hours |
| OM Opt. (Diffusion Model) | No | 1e4 | 50 min |
Table 4 presents results from Ramakrishnan et al. (2025) comparing the Nudged Elastic Band (NEB) method with their Implicit Neural Representation (INR) approach for MEP discovery. Both our AdaPath approach and their INR method utilize lightweight, preexisting molecular dynamics force fields, though we employ different such force fields. Their approach requires a more complex loss function that evaluates both energy and forces along with path parameterization and its gradient (path velocity). However, it needs fewer force field evaluations than our method. It is important to note that the ”Max Energy” reported in Table 4 is measured in electron volts (eV) and appears to be adjusted for ground state energy reference, which makes direct energy comparisons with our method challenging. The difference in force fields used further complicates any direct energy value comparisons between methods.
These tables corroborate our findings in the Doob’s Lagrangian comparison: methods that optimize for a single MEP consistently require orders of magnitude fewer force field evaluations than approaches that characterize the complete transition path ensemble.
E.1 Connection between Doob’s Lagrangian and MEP Optimization